OpenStudy (anonymous):

write an equation of the line containing the given point and parallel to the given line. (-5,3); 9x=2y+3

OpenStudy (anonymous):

parallel lines have same slope so y=mx+b then use point and slope to find b

OpenStudy (anonymous):

9x-3=2y y=9/2 x - 3/2 so slope is 9/2

OpenStudy (anonymous):

plug in 9/2 and -5 for x and 3 for y and solve for b

OpenStudy (anonymous):

are you following me

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

after i plug it in than what is next

OpenStudy (anonymous):

what did you get for b

OpenStudy (anonymous):

b=51/2 so 3=-45/2 + b add -45/2 to both sides and change 3 to 6/2 so you can add 45/2 y=9/2 x +51/2

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

welcome...fan me

OpenStudy (anonymous):

I'm new to the site how do I do that

OpenStudy (anonymous):

click "Become a fan" in blue beside my name

OpenStudy (anonymous):

I'm new too so welcome

OpenStudy (anonymous):

do you have any more questions

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(7,9); 2x+y=3

OpenStudy (anonymous):

checking to see if it is a solution?

OpenStudy (anonymous):

2(7) + 9 = 3 14+ 9 = 3 NOOOO so not a solution

OpenStudy (anonymous):

write an equation of the line containing the given point and parallel to the given line (5,-5); 4x+3y=7

OpenStudy (anonymous):

transform to y=mx+b by subtracting 4x from bot h sides then divide both sides by 3 to get y =

OpenStudy (anonymous):

do that first and tell me what you get

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

would it b 12

OpenStudy (anonymous):

y=-4/3x + 7/3 now plug in (-5,5) and slope = -4/3 into y=mx+b and solve for b

OpenStudy (anonymous):

(5, -5) sorry

OpenStudy (anonymous):

-5=-4/3(5) + b

OpenStudy (anonymous):

(5, -5) sorry

OpenStudy (anonymous):

-5=(-4/3)(5)+b

OpenStudy (anonymous):

-5=-20/3 +b add 20/3 so 5 which is 15/3+20/3=35/3

OpenStudy (anonymous):

-15/3 + 20/3 = 5/3

OpenStudy (anonymous):

b=5/3 same slope of -4/3 so y=-4/3x +5/3

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

no

OpenStudy (anonymous):

what is wrong

OpenStudy (anonymous):

do you need me to put it all together in one post

OpenStudy (anonymous):

please

OpenStudy (anonymous):

4x+3y=7 solve for y -4x -4x 3y=7-4x or 3y=-4x+7 ---------- 3 y=-4/3 x + 7/3 so slope is -4/3 parallel lines have same slope so use m=-4/3 and (5, -5) and plug into slope-intercept form to find "b"-y intercept y=mx+b -5=-4/3(5)+b -5=-20/3 + b +20/3 +20/3 -15/3 + 20/3 = b 5/3= b so y=-4/3x + 5/3

OpenStudy (anonymous):

got it now?